OpenStudy (anonymous):

Ignore the last one Log[d]12= 6

jimthompson5910 (jim_thompson5910):

\[\Large \log_{d}(12) = 6\] \[\Large d^6 = 12\] \[\Large \sqrt[6]{d^6} = \sqrt[6]{12}\] \[\Large d \approx 1.5130857 \]